Honiew Cooper (1.5-Mile) Run Test Height (cm): 180 Weight (kg): 85.0 Optional: Heart rate (bpm): 196 miles Initials: AA Age(y): 20 Run time (min:sec): 11 laps Optional: Distance run in 12 min: Fitness category (Table 12.3124) VO-max (Table 12.1/12.2) Evaluation/ comments: ml kg min George (1-Mile) Jog Test Gender: F Initials: BB Age (y) 20 Gender (M-1,F-О) Jog time (min:sec) 10 : VO,max 100.5+(8.344(0.1636(1.438 Height (cm): 160 Weight (kg): 65.0 45 Jog time (min) HR (bpm) 171 -(0.1928 Heart rate Body wi Jog time Gender Vo,max (mkkg-min) Evaluation/ comments: Fitness category (Table 12.5) Rockport (1-Mile) Walk Test (Modified for 18-29 y) Gender: F Initials: CC Age () 20 Height (cm): 165 Weight (kg): Gender (M-1,F-o) walk time (min:sec) 16:14 Walk time (min) VO2max 88.768 + (8.892 * 65.0 HR (bpm) 131 )-(0.21 09-(1.4537-(0.1194 Gender Body wt Walk time Heart rate Vo,max (ml-kg min) Fitness category (Table 12.7) Evaluation/comments: CHAPTER 12 Aerobic Running, Jogging, and Walking 1
Answer #1

Cooper (1.5 mile) run test:

Gender- Male Age- 20 Height- 180 cm

weight- 85 Kg Heart rate- 196 bpm

Distance run in 12 minutes: 6 laps = 1.5 miles

VO2max= 35.97 * miles = 35.97 * 1.5 = 42.67 ml.kg-1.min-1

Since the fitness category chart isn't mentioned in the question, I am using the table, taken from TopEndSports, that shows normative data for VO2max in different population groups.

Fitness category: Average.

Evaluation/ Comments: The score is within average levels, which is good for a regular runner. He can increase his VO2 max by training regularly and doing plenty of high intensity interval training workouts.

George (1 mile) jog test:

Gender: Female Age: 20 yrs. Height: 160 cm   

Weight: 65 kg

Jog time: 10 min 45 secs

Jog time in minutes: 10.75 minutes Heart rate: 171 bpm

VO2max: 100.5 + (8.344 * gender) - (0.1636* Weight) - (1.436* jog time) - (0.1928* Heart rate)

= 100.5 + (8.344* 0) - (0.1636 * 65) - (1.436* 10.75) - (0.1928* 171)

= 41.46 ml.Kg-1.min-1

Fitness category: Average.

Evaluation/ Comments: The score is within average levels, which is good for a regular runner. She can increase her VO2 max by training regularly and doing plenty of high intensity interval training workouts.

Rockport (1 mile) walk test:

Gender: Female Age: 20 yrs. Height: 165 cm   

Weight: 65 kg

Walk time time: 16 min 14 secs

Walk time in minutes: 16.23 minutes Heart rate: 131 bpm

VO2max: 88.768 + (8.892* gender) - (0.2109* Weight) - (1.4537* walk time) - (0.1194* Heart rate)

= 88.768 + (8.892* 0) - (0.2109* 65) - (1.4537*16.23 ) - (0.1194* 131)

= 35.83 ml.kg-1.min-1

Fitness category: Below average.

Evaluation/ Comments: She can increase her VO2 max by training regularly and doing plenty of high intensity interval training workouts.While a good VO2max can certainly contribute to one's success, particularly with endurance sports, there are other factors that arguably play a larger role, including skills training, psychological preparation, lactate threshold training, and nutrition.